# Detecting Functional Bugs in Smart Contracts through LLM-Powered and Bug-Oriented Composite Analysis

### v1

@@ 3.2. Invariant Checker Generation
Although SmartInv (Wang et al., 2024) utilizes a pre-trained LLaMA-7B model to infer invariants from smart contracts automatically, its generated invariants suffer from low-quality issues due to the following reasons. First, the model’s outputs frequently include hallucinations that significantly compromise the reliability of the generated invariants. Second, the presence of existing logical judgments in the code can obscure key logical information related to potential bugs, impacting the model’s ability to accurately generate relevant invariants. Third, the small parameter size of the LLaMA-7B model necessitates a well-designed dataset for effective fine-tuning to improve its performance in generating precise and useful invariants.
@@
Table 2. Characteristics of critical variables and principal statements for each bug subcategory, along with the associated invariant checkers.
Bug Subcategory Key Variables (V) & Statements (S) Invariant Checker
AMM Price Oracle Manipulation V1: calculated price of the LP token PriceChange_Checker(V1)
Non-AMM Price Oracle Manipulation V1: tokens spent by the user during the swap process ExchangeRate_Checker(V1,V2)
V2: the amount of tokens received by the user after the swap
Approval Not Clear V1: sending address for token transfers TokenChange_Checker(V1,V2,V3)
V2: authorization array for token transfers
V3: the amount of tokens being transferred
Unauthorized Transfer V1: sending address for token transfers TokenChange_Checker(V1,V2,V3)
V2: authorization array for token transfers
V3: the amount of tokens being transferred
Wrong Checkpoint Order S1: the first to call Checkpoint StatementOrder_Checker(S1,S2)
S2: the first to perform a calculation or update on the tokens
Risky First Deposit V1: the total supply of tokens ShareSafety_Checker(V1,V2,V3)
V2: the share that a user can obtain
V3: tokens spent by the user
Wrong Interest Rate Order S1: the first to update the interest rate StatementOrder_Checker(S1,S2)
S2: the first to use the interest rate to calculate or update the
token balance, rewards, or shares
Improper Handling of Deposit Fee V1: tokens spent by the user without any deductions TokenChange_Checker(V1,V2,V3)
V2: the net amount added to the user’s balance after deducting
the deposit fee
V3: the amount of deposit fee
Wrong Implementation of Amount Lock
V1: the array where the first element is used to store the user’s
original locked amount
TokenChange_Checker(V1,V2,V3)
V2: the amount by which the user’s locked funds have increased
or decreased
V3: stores the new locked amount after increasing or decreasing
based on the isIncrease flag
Vote Manipulation V1: contract variable for voting tokens StateChange_Checker(V1,V2,S1)
V2: the number of votes cast by the user in this round
S1: the first to execute a voting action
@@
In response to this challenge, we propose a hierarchical matching method that capitalizes on the capabilities of GPT for the extraction of critical variables and principal statements. Our approach unfolds in two primary steps. First, we perform an in-depth analysis of each bug category to pinpoint critical variables and principal statements that are vital for comprehending their logic. For instance, in the scenario of an AMM price oracle manipulation bug, a critical variable is the one that holds the calculated price of the LP token. We present the characteristics of critical variables and principal statements for each bug subcategory in Table 2. Second, leveraging the identified elements, we formulate specific prompts and input them into GPT to query each vulnerable function as identified by the LLM-driven multi-perspective analysis module. Figure 6 illustrates the prompt template designed to extract critical variables and principal statements from code snippets potentially vulnerable to AMM price oracle manipulation. For different bug categories, we can substitute [%Critical Variable%] and [%Principal Statement%] in the template with the predefined elements specific to those bugs. With these tailored prompts, GPT is able to precisely pinpoint the relevant critical variables and principal statements, enhancing the accuracy of our bug detection process.
@@
Figure 6. Prompt template for extracting critical variables from code snippets potentially vulnerable to AMM price oracle manipulation.
@@
Template-based Checker Generation. While we have extracted critical variables and principal statements from smart contracts, we currently lack an efficient method to convert these into usable invariants for bug detection. To fill this gap, we propose a template-based checker generation method. Specifically, for each type of bug, we design a corresponding invariant checker template that targets the vulnerable features of different bugs, as detailed in Table 2. We mainly design 6 types of checkers: PriceChange_Checker, ExchangeRate_Checker, TokenChange_Checker, StatementOrder_Checker, ShareSafety_Checker, and StateChange_Checker. Each checker includes a conditional statement that assesses the contract state. For instance, as shown in Figure 7, the PriceChange_Checker for AMM price oracle manipulation determines whether the variable (i.e., calculated price of the LP token) falls below $90\%$ or exceeds $110\%$ of the old price. Such a condition indicates that the price change surpasses the ±10% threshold, typically signaling abnormal price fluctuations and triggering an alert for an existing AMM price oracle manipulation bug. These checkers could be dynamically combined to comprehensively cover a broad range of functional bugs, ensuring scalability.
Figure 7. The checker for AMM price oracle manipulation.
3.3. Bug-Oriented Analysis Engine
Though we have identified potentially vulnerable functions in smart contracts and generated the corresponding invariant checkers, we still lack a practical method to leverage them for functional bug detection. Existing works mainly rely on static analysis tools to validate rule violations or invariant breaches to detect functional bugs. For instance, GPTScan leveraged Falcon (met, 2024), a static analysis tool built upon Slither (Feist et al., 2019), to confirm the existence of the candidate vulnerable functions. Similarly, SMARTINV employed a customized static checker based on VERISOL (Wang et al., 2019a) to detect violations of inferred invariants. Nevertheless, these tools often suffer from false positives and cannot fully simulate contract behavior in real-world environments, potentially missing runtime-specific issues, leading to false negatives. To tackle the above problem, we implement a bug-oriented analysis engine with the following designs.
@@ 3.3. Bug-Oriented Analysis Engine
Bug-Oriented Fuzzing. After inserting invariant checkers into smart contracts, we then employ fuzzing to address the aforementioned limitations in static analysis methods. Traditional dynamic analysis techniques often explore programs randomly without a specific target, typically identifying bugs through program crashes. However, functional bugs in smart contracts do not generally lead to crashes and are less likely to be detected through random testing. To mitigate this, we implement a bug-oriented fuzzing approach that directs the fuzzer to specifically target functions with inserted invariant checkers, rather than performing random exploration. Our method builds upon ItyFuzz (Shou et al., 2023), a state-of-the-art hybrid fuzzer for smart contracts that integrates symbolic execution with fuzzing. ItyFuzz has shown great performance in finding implementation bugs but falls short in detecting functional bugs based on our evaluation in Section 4.4. Therefore, to bridge the gap, we customize ItyFuzz by leveraging its functionality that allows verification of custom invariants. This customization involves inserting a specific bug alert function into each invariant checker, which allows us to direct the fuzzer to activate these targeted checkers rather than engaging in random exploration. Our method significantly improves the efficiency and effectiveness of the fuzzing process.
0:   $C$: smart contract code $B$: potential functional bug subcategory 0:   $R$: the set of detected functional bug $Q$: the set of results requiring further inspection 1:  for each $B_{i}\in B$ do 2:     $v,s=Extract\_Variables\_and\_Statements(B_{i},C)$ 3:     $inv=Generating\_invariants(v,s)$ 4:     $CwI=Invariant\_Checker\_Insertion(inv,B_{i},C)$ 5:     $C_{s}\leftarrow initial\_infant\_state\_corpus$ 6:     $C_{ts}\leftarrow initial\_transaction\_and\_state\_pair\_corpus$ 7:     while $t\neq\varnothing$ do 8:        $s_{mut},t_{mut}\leftarrow input\_mutation(C_{ts},C_{s})$ 9:        $f,s^{\prime},violation\leftarrow Execution(s_{mut},t_{mut},CwI)$ 10:        if  $violation\ is\ found$  then 11:           $R=R.append(inv,B_{i})$ 12:        end if 13:        $update\_corpus(f,s^{\prime},s_{mut},t_{mut})$ 14:     end while 15:     if  $violation\ is\ not\ found$  then 16:        $Q=Q.append(inv,B_{i})$ 17:     end if 118:  end for 19:  return  $R$, $Q$
Algorithm 2 Functional Bug Detection
3.4. Functional Bug Detection
The functional bug detection module aims to perform comprehensive bug detection on smart contracts. Algorithm 2 outlines the procedure for applying this methodology. Specifically, in this module, we first compile the target smart contracts with strategically inserted invariant checkers. Next, we perform fuzzing on these smart contracts. When the fuzzer activates an invariant checker and detects a discrepancy, it signals the presence of a bug. According to the results, we then generate a comprehensive bug report that documents all identified functional bugs within the smart contracts. This report serves as a crucial tool for developers to understand and rectify bugs in their smart contracts.
4. System Evaluation
@@
RQ1: How does the dual-agent prompt engineering strategy contribute to the performance of PromFuzz? (Section 4.2)
RQ2: What is PromFuzz’s accuracy in generating invariant checkers? (Section 4.3)
@@
RQ3: How effective is PromFuzz at detecting functional bugs in smart contracts? (Section 4.4)
@@
RQ4: Is PromFuzz capable of discovering zero-day functional bugs in real-world smart contracts? (Section 4.5)
4.1. Dataset
@@
To ensure a comprehensive assessment of PromFuzz’s effectiveness, we create the following 6 datasets covering a diverse range of functional bug categories. These datasets are sourced from DeFiHackLabs (def, 2024), Web3bugs (Zhang et al., 2023), GPTScan (Sun et al., 2024), on-chain smart contracts and real-world DeFi projects. We selectively use these datasets, excluding projects lacking necessary external files or configuration information for compilation, as PromFuzz’s dynamic analysis necessitates complete setups. Furthermore, datasets like SmartBugs (Durieux et al., 2020) and VeriSmart (So et al., 2020), which lack functional bugs, are excluded from our evaluation.
Bug-free smart contract dataset ($D_{BF-SC}$) includes well-audited and widely used smart contracts that are free from any known functional bugs. These contracts are sourced from GPTScan (Sun et al., 2024) and serve as a benchmark to assess the false positive rate of PromFuzz. Specifically, $D_{BF-SC}$ consists of $303$ open-source contract projects from $6$ mainstream Ethereum-compatible chains.
@@
Exploited smart contract dataset ($D_{Exploit-SC}$) includes smart contracts that have been subjected to real-world attacks in the past, each causing significant monetary losses. $D_{Exploit-SC}$ is sourced from DeFiHackLabs (def, 2024) and contains $7$ real-world attack cases that we have manually verified as being caused by functional bugs. Among these, $4$ cases are due to price oracle manipulation, $2$ are due to unauthorized behavior, and $1$ is due to incorrect control mechanism.
@@
Bug-verified smart contract dataset ($D_{BV-SC}$) includes smart contracts audited through platforms like Code4rena (Cod, 2024) and Immunefi (Imm, 2024), with confirmed high-risk functional bugs by the developers. $D_{BV-SC}$ is sourced from Web3bugs (Zhang et al., 2023), which includes $10$ functional bugs across $36$ smart contract files. Among these, $5$ bugs result from insecure calculating logic, $2$ are due to incorrect control mechanisms, $2$ are due to unauthorized behavior, and $1$ is due to price oracle manipulation.
Synthetic bug-injected smart contract dataset ($D_{Synth-SC}$) includes smart contracts with manually injected functional bugs. The purpose of this dataset is to enlarge $D_{Exploit-SC}$ and $D_{BV-SC}$, providing a comprehensive evaluation of PromFuzz. We create $D_{Synth-SC}$ by manually injecting functional bugs into well-performing, bug-free projects on the Ethereum-compatible chain, which includes $6$ manually injected functional bugs across $44$ smart contract files. Among these, $4$ bugs are due to unauthorized behavior and $2$ are due to incorrect control mechanisms.
Symbol-extracted smart contract dataset ($D_{SE-SC}$) includes smart contact functions with manually extracted and verified critical variables and principal statements. Specifically, $D_{Extracted-SC}$ includes $55$ extracted critical variables and principal statements across $23$ smart contract functions.
@@
Real-world DeFi project dataset ($D_{Real-DeFi}$) comprises DeFi projects currently undergoing public audits through various platforms. Specifically, $D_{Real-DeFi}$ includes $6$ such DeFi projects.
Table 3. False positive rate evaluation.
Tool $D_{BF-SC}$
FPR (Function-Level) FPR (File-Level)
PromFuzz 6.31% 15.38%
GPTScan 7.16% 15.75%
SMARTINV N/A 72%
Table 4. Comparison of Dual-agent and Auditor-agent.
Tool $D_{Exploit-SC}$ $D_{BV-SC}$ $D_{Synth-SC}$ Overall
TP FP FN TP FP FN TP FP FN TP FP FN Recall F1-Score
Dual-agent 6 16 1 9 12 1 6 6 0 21 34 2 91.30% 53.85%
Auditor-agent 5 9 2 7 10 3 3 5 3 15 24 8 65.22% 48.39%
4.2. Dual-Agent Architecture Evaluation
This subsection addresses RQ1 by assessing the effectiveness of the dual-agent architecture of PromFuzz. We begin with evaluating the false positive rate of the LLM-driven multi-perspective module in PromFuzz. While PromFuzz generally mitigates false positives during the fuzzing phase, the occurrence of additional false positives requires PromFuzz to expend more resources on verifying non-existent bugs. These unnecessary validation operations will incur additional costs in generating invariant checkers and consume substantial time during the fuzzing phase. To this end, we compare PromFuzz with two established tools: GPTScan and SMARTINV. GPTScan integrates GPT with static analysis to detect functional bugs in smart contracts. SMARTINV leverages a fine-tuned LLM to infer invariants from smart contracts and performs further analysis based on a static tool. In this experiment, PromFuzz is exclusively compared with the LLM modules of GPTScan and SMARTINV. We set up both tools according to the methodologies described in their respective papers and then perform them on $D_{BF-SC}$. Both GPTScan and PromFuzz utilize GPT-4-turbo as their foundational model.
@@ 4.2. Dual-Agent Architecture Evaluation
Next, we conduct an ablation study to evaluate the performance differences between the dual-agent based PromFuzz and a version that utilizes only the Auditor Agent. To achieve this goal, we construct two new PromFuzz, the first one is PromFuzz + Dual-agent, and the second one is PromFuzz + Auditor-agent. This comparison is conducted using three datasets: $D_{Exploit-SC}$, $D_{BV-SC}$, and $D_{Synth-SC}$.
False Positive Rate Evaluation. Both PromFuzz and GPTScan can analyze whether a function contains functional bugs, operating at the function-level, while SMARTINV is limited to file-level analysis. For a fair comparison, we conduct evaluations at both the function and file levels. As illustrated in Table 3, PromFuzz has a slight advantage with false positive rates of $6.31\%$ and $15.38\%$, respectively, than GPTScan. Despite strictly adhering to the guidelines for SMARTINV and utilizing its publicly released code, SMARTINV records a significantly higher false positive rate of $72\%$ at file-level.
We further analyze the false positives of these tools. For SMARTINV, the primary focus of its LLaMA-7B fine-tuning strategy and prompt design is on generating critical invariants. However, this approach lacks comprehensive reasoning capabilities necessary for effective bug detection. This deficiency in the analytical strategy significantly contributes to the tool’s notably high false positive rate. For GPTScan, the primary reason for false positives arises from its single-agent design, which can lead to the incorrect identification of correct behaviors as functional bugs, such as those involving insecure calculating logic. Although PromFuzz has similar reasons, its innovative dual-agent design effectively mitigates a greater number of false positives compared to single-agent systems.
@@
Ablation Study. As shown in Table 4, Dual-agent shows substantial performance improvements over the single-agent version. Specifically, on the $D_{Exploit-SC}$ dataset, Dual-agent identifies 6 true positives, 16 false positives, and only 1 false negative. On the $D_{BV-SC}$ dataset, it detects 9 true positives, 12 false positives, and 1 false negative. Notably, on the $D_{Synth-SC}$ dataset, Dual-agent identifies 6 true positives, 6 false positives, and no false negatives. Overall, Dual-agent achieves a recall of 91.30% and an F1-score of 53.85%, identifying 21 true positives and 34 false positives across all datasets, with only 2 false negatives. In contrast, Auditor-agent detects fewer true positives and has more false negatives, achieving a recall of 65.22% and an F1-score of 48.39%.
@@
By adopting our dual-agent prompt engineering strategy, PromFuzz + Dual-agent significantly improves true positives and eliminates false negatives. Although PromFuzz + Dual-agent exhibits a higher number of false positives, these are manageable as they can be effectively filtered out during the subsequent fuzzing phase.
@@
We further analyze the false positives and false negatives by PromFuzz + Dual-agent and PromFuzz + Auditor-agent. For PromFuzz + Dual-agent, though the Attacker Agent increases true positives by considering multiple angles, it still has limitations and can lead to over-generation of false positives. For PromFuzz + Auditor-agent, auditor’s perspective, which focuses solely from an auditing standpoint, can result in underreporting, filtering out some high-risk vulnerabilities. The limitations of auditor prompt rules, which, like those of the attacker, cannot cover all bugs, resulting in some being missed. Besides, for both PromFuzz + Dual-agent and PromFuzz + Auditor-agent, GPT’s understanding is still insufficient, leading to cases of prompt misinterpretation.
4.3. Invariant Checker Generation Accuracy

### v2

@@ III-B Invariant Checker Generation
In response to this challenge, we propose a hierarchical matching method that capitalizes on the capabilities of GPT for the extraction of critical variables and principal statements. Our approach unfolds in two primary steps. First, we perform an in-depth analysis of each bug category to pinpoint critical variables and principal statements that are vital for comprehending their logic. For instance, in the scenario of an AMM price oracle manipulation bug, a critical variable is the one that holds the calculated price of the LP token. We present the characteristics of critical variables and principal statements for each bug subcategory in our supplementary material(§II). Second, leveraging the identified elements, we formulate specific prompts and input them into GPT to query each vulnerable function as identified by the LLM-driven multi-perspective analysis module. Figure 5 illustrates the prompt template designed to extract critical variables and principal statements from code snippets potentially vulnerable to AMM price oracle manipulation. For different bug categories, we can substitute [%Critical Variable%] and [%Principal Statement%] in the template with the predefined elements specific to those bugs. With these tailored prompts, GPT is able to precisely pinpoint the relevant critical variables and principal statements, enhancing the accuracy of our bug detection process.
@@
Figure 5: Prompt template for extracting critical variables from code snippets potentially vulnerable to AMM price oracle manipulation.
@@
Template-based Checker Generation. While we have extracted critical variables and principal statements from smart contracts, we currently lack an efficient method to convert these into usable invariants for bug detection. To fill this gap, we propose a template-based checker generation method. Specifically, we mainly design $6$ types of checkers: PriceChange_Checker, ExchangeRate_Checker, TokenChange_Checker, StatementOrder_Checker, ShareSafety_Checker, and StateChange_Checker. For a detailed invariant checker templates, please see our supplementary material(§III). Each checker includes the conditional statements that assesses the contract state. For instance, as shown in Figure 6, the PriceChange_Checker for AMM price oracle manipulation determines whether the variable (i.e., calculated price of the LP token) falls below $90\%$ or exceeds $110\%$ of the old price. Such a condition indicates that the price change surpasses the ±10% threshold, typically signaling abnormal price fluctuations and triggering an alert for an existing AMM price oracle manipulation bug. These checkers could be dynamically combined to comprehensively cover a broad range of functional bugs, ensuring scalability.
Figure 6: The checker for AMM price oracle manipulation.
0:
$C$: smart contract code
$B$: potential functional bug subcategory
10:
$R$: the set of detected functional bugs
1ex
1:  for each $B_{i}\in B$ do
2:   $v,s=Extract\_Variables\_and\_Statements(B_{i},C)$
3:   inv=Generating_invariants(v,s)inv=Generating\_invariants(v,s)
4:   $CwI=Invariant\_Checker\_Insertion(inv,B_{i},C)$
5:   $C_{s}\leftarrow initial\_infant\_state\_corpus$
6:   $C_{ts}\leftarrow initial\_transaction\_and\_state\_pair\_corpus$
7:   while $t\neq\varnothing$ do
8:    $s_{mut},t_{mut}\leftarrow input\_mutation(C_{ts},C_{s})$
9:    $f,s^{\prime},violation\leftarrow Execution(s_{mut},t_{mut},CwI)$
10:    if violationisfoundviolation\ is\ found then
11:     $R=R.append(inv,B_{i})$
12:    end if
13:    $update\_corpus(f,s^{\prime},s_{mut},t_{mut})$
14:   end while
15:  end for
16:  return $R$
Algorithm 1 Functional Bug Detection
III-C Bug-Oriented Analysis Engine
Though we have identified potentially vulnerable functions in smart contracts and generated the corresponding invariant checkers, we still lack a practical method to leverage them for functional bug detection. To tackle the above problem, we implement a bug-oriented analysis engine with the following designs.
@@ III-C Bug-Oriented Analysis Engine
Bug-Oriented Fuzzing. After inserting invariant checkers into smart contracts, we then employ fuzzing to address the aforementioned limitations in static analysis methods. Traditional dynamic analysis techniques often explore programs randomly without a specific target, typically identifying bugs through program crashes. However, functional bugs in smart contracts do not generally lead to crashes and are less likely to be detected through random testing. To mitigate this, we implement a bug-oriented fuzzing approach that directs the fuzzer to specifically target functions with inserted invariant checkers, rather than performing random exploration. Our method involves inserting a specific bug alert function into each invariant checker, which allows us to direct the fuzzer to activate these targeted checkers rather than engaging in random exploration. Our method significantly improves the efficiency and effectiveness of the fuzzing process.
III-D Functional Bug Detection
The functional bug detection module aims to perform comprehensive bug detection on smart contracts. Algorithm 1 outlines the procedure for applying this methodology. Specifically, in this module, we first compile the target smart contracts with strategically inserted invariant checkers (line 2-4). Next, we perform fuzzing on these smart contracts (line 5-14). When the fuzzer activates an invariant checker and detects a discrepancy, it signals the presence of a bug (line 10-12). According to the results, we then generate a comprehensive bug report that documents all identified functional bugs within the smart contracts.
IV System Evaluation
@@
RQ1: How does the dual-agent prompt engineering strategy contribute to the performance of PromFuzz? (Section IV-B)
RQ2: What is PromFuzz’s accuracy in generating invariant checkers? (Section IV-C)
@@
RQ3: How effective is PromFuzz at detecting functional bugs in smart contracts? (Section IV-D)
@@
RQ4: Is PromFuzz capable of discovering zero-day functional bugs in real-world smart contracts? (Section IV-E)
IV-A Dataset
@@
To ensure a comprehensive assessment of PromFuzz’s effectiveness, we systematically collect $459$ smart contracts from various sources, including DeFiHackLabs [6], Web3bugs [45], SmartBugs [12], VeriSmart [32], on-chain smart contracts, and real-world DeFi projects. After further analysis, we remove $198$ contracts that either lack functional bugs or are missing essential external files or configuration information required for compilation. Ultimately, we create $5$ datasets comprising $261$ contracts, covering a broad spectrum of functional bug categories and evaluation perspectives. A detailed description of these datasets is provided in Table I.
TABLE I: The evaluation datasets.
Dataset Description RQ
$D_{Exploit-SC}$ 7 real-world attack cases with 47 contracts RQ1,RQ3
$D_{BV-SC}$ 10 real-world audit cases with 36 contracts RQ1,RQ3
$D_{Synth-SC}$ 6 real-world synthetic cases with 44 contracts RQ1,RQ3
$D_{SE-SC}$ 26 contracts with 59 extracted symbols RQ2
$D_{Real-DeFi}$ 10 real-world DeFi projects with 108 contracts RQ4
@@
Exploited smart contract dataset ($D_{Exploit-SC}$) includes smart contracts that have been subjected to real-world attacks in the past, each causing significant monetary losses. $D_{Exploit-SC}$ includes $7$ real-world attack cases, associated with $47$ contracts, all of which we have manually verified to be caused by functional bugs.
@@
Bug-verified smart contract dataset ($D_{BV-SC}$) includes smart contracts audited through platforms like Code4rena [2] and Immunefi [4]. $D_{BV-SC}$ includes $10$ real-world audit cases, associated with $36$ contracts, all of which have been confirmed by the developers to contain high-risk functional bugs.
Synthetic smart contract dataset ($D_{Synth-SC}$) consists of smart contracts where functional issues have been deliberately introduced. The purpose of this dataset is to expand $D_{Exploit-SC}$ and $D_{BV-SC}$, enabling a more comprehensive evaluation of PromFuzz. $D_{Synth-SC}$ includes $6$ synthetic cases across $44$ smart contracts. We construct $D_{Synth-SC}$ with the following steps. First, we collect smart contracts that are well-designed, well-audited, and widely deployed on mainstream blockchains, which reasonably indicates a low likelihood of major bugs. Then, two of our authors, who are experienced in smart contract development, perform controlled bug injection. One author implements the injection based on bug patterns summarized from real-world attack incidents, while the other conducts independent verification.
Symbol-extracted smart contract dataset ($D_{SE-SC}$) includes smart contract functions with manually extracted and verified critical variables and principal statements. Specifically, $D_{SE-SC}$ includes $59$ extracted critical variables and principal statements across $26$ smart contract functions. In the construction of $D_{SE-SC}$, we assign two of our authors with expertise in smart contract security. The first author conducted a manual analysis of each smart contract, identifying critical variables and principal statements relevant to invariant checker generation and functional bug detection. The second author performed a secondary verification to ensure the accuracy and reliability of the extracted information.
@@
Real-world DeFi project dataset ($D_{Real-DeFi}$) comprises DeFi projects currently undergoing public audits through various platforms. Specifically, $D_{Real-DeFi}$ includes $10$ such DeFi projects with $108$ smart contracts.
TABLE II: Comparison of Dual-agent and Auditor-agent.
Architecture Overall
TP FP FN Recall F1-Score
Dual-agent 21 34 2 91.30% 53.85%
Auditor-agent 15 24 8 65.22% 48.39%
IV-B Dual-Agent Architecture Evaluation
@@
This subsection addresses the effectiveness of the dual-agent architecture of PromFuzz. First, we conduct an ablation study to evaluate the performance differences between the Dual-agent based PromFuzz and a version that utilizes only the Auditor Agent. Second, we compare Dual-agent with two mainstream reasoning models: DeepSeek-R1 and OpenAI-o3. The above comparisons are conducted using three datasets: $D_{Exploit-SC}$, $D_{BV-SC}$, and $D_{Synth-SC}$.
@@
Ablation Study. As shown in Table II, Dual-agent architecture shows substantial performance improvements over the Auditor-agent architecture. Specifically, Dual-agent architecture achieves a recall of 91.30% and an F1-score of 53.85%, identifying 21 true positives and 34 false positives across $D_{Exploit-SC}$, $D_{BV-SC}$, and $D_{Synth-SC}$ datasets, with only 2 false negatives. In contrast, Auditor-agent architecture detects fewer true positives and has more false negatives, achieving a recall of 65.22% and an F1-score of 48.39%.
@@
By adopting our dual-agent prompt engineering strategy, Dual-agent architecture significantly improves true positives and eliminates false negatives. Although Dual-agent architecture exhibits a higher number of false positives, these are manageable as they can be effectively filtered out during the subsequent fuzzing phase.
@@
We further analyze the false positives and false negatives by Dual-agent architecture and Auditor-agent architecture. For Dual-agent architecture, though the Attacker Agent increases true positives by considering multiple angles, it still has limitations and can lead to over-generation of false positives. For Auditor-agent architecture, auditor’s perspective, which focuses solely from an auditing standpoint, can result in underreporting, filtering out some high-risk vulnerabilities. The limitations of auditor prompt rules, which, like those of the attacker, cannot cover all bugs, resulting in some being missed. Besides, for both Dual-agent architecture and Auditor-agent architecture, GPT’s understanding is still insufficient, leading to cases of prompt misinterpretation.
Comparison with reasoning models. As shown in Table III, Dual-agent architecture outperforms mainstream reasoning model architectures. Specifically, Dual-agent architecture ahieves a recall of 91.30% and an F1-score of 53.85%. OpenAI-o3 architecture achieves a recall of 52.17% and an F1-score of 32.43%. DeepSeek-R1 architecture achieves a recall of 47.83% and an F1-score of 39.29%.
We further analyze the false positives and false negatives by mainstream reasoning models. In OpenAI-o3 and DeepSeek-R1, the decomposition of the detection task is determined autonomously by the reasoning model. Such task planning can introduce a degree of randomness, potentially guiding the LLM toward incorrect reasoning paths and resulting in false positives and false negatives. In contrast, Dual-agent architecture designs its reasoning workflow based on strategies observed in real-world auditors and attackers, enabling the LLM to follow a more structured and reliable analysis process.
TABLE III: Comparison of Dual-agent and reasoning models.
Architecture Overall
TP FP FN Recall F1-Score
Dual-agent 21 34 2 91.30% 53.85%
OpenAI-o3 12 39 11 52.17% 32.43%
DeepSeek-R1 11 22 12 47.83% 39.29%
IV-C Invariant Checker Generation Accuracy